LUCI AO

0:26
  • Waiting for them to do some MODS observations.
4:34
  • Ready to start, cooling the OCAMS. Seeing looks really good.
  • Bin4 297 Hz 16.5 mag
  • I think that bin 4 won't work for seeing estimation, but let's try it.
  • 20240301_045117
  • It works, just not very well.
  • 20240301_045549
  • 20240301_045638
  • 20240301_045733
  • 20240301_045812
5:49
  • New Target
  • Bin3 398 Hz Mag 15.4
  • 20240301_055025
  • Optical gain is off! That's the issue.
  • I don't trust it in bin3, and the seeing is really stable, so It's staying off.
6:34
  • New Target
  • Bin3 375 Hz 15.5 mag
  • Going to try GOPT. It looks like it should work just fine.
  • 20240301_063642
  • 20240301_063852
  • No DIMM yet
  • 20240301_064053
  • DIMM's back
  • 20240301_064231
  • 20240301_064403
  • These datasets are giving me about 1/2 of what DIMM value is.
  • I guess that we aren't going fast enough for the ASM modes to track the actual seeing, or something.
7:14
  • New Target
  • Bin3 372Hz Mag 15.5
  • Not much point in taking datasets as GOPT isn't all that stable in bin3
8:00
  • Last Target
  • Bin4 266Hz 16.5 mag
8:40
  • Done!

-- GregoryTaylor - 01 Mar 2024
Topic revision: r2 - 01 Mar 2024, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback